# A Near-optimal, Scalable and Corruption-tolerant Framework for Stochastic Bandits: From Single-Agent to Multi-Agent and Beyond

### v1

@@ 4.2 $d$-Set Semi-Bandits
1
2Initialization: Set the initial round $T_{0}=0$, $\Delta_{k}^{0}=1$, and $r_{k}^{0}=0$ for all $k\in[K]$.
3for epochs $m=1,2,\cdots$ do
4       Set $\zeta_{m}\leftarrow(m+4)2^{2(m+4)}\ln(K)$, and $\delta_{m}\leftarrow 1/(K\zeta_{m})$
5      Set $\lambda_{m}\leftarrow 2^{8}\ln{\left(4K/\delta_{m}\right)}$ and $\beta_{m}\leftarrow\delta_{m}/K$.
6      Set $n_{k}^{m}=\lambda_{m}(\Delta_{k}^{m-1})^{-2}$ for all arms $k\in[K]$.
7      Set $N_{m}\leftarrow\lceil K\lambda_{m}2^{2(m-1)}/d\rceil$ and $T_{m}\leftarrow T_{m-1}+N_{m}$.
8      Select the arm subsets ${\mathcal{K}}_{m}=\mathop{\arg\max}_{k\in[K]}r_{k}^{m-1}$ with $|{\mathcal{K}}_{m}|=d$.
9      Set $\widetilde{n}_{k}^{m}=\begin{cases}n_{k}^{m}&k\not\in{\mathcal{K}}_{m}\\ N_{m}-\sum_{k\not\in k_{m}}n_{k}^{m}/d&k\in{\mathcal{K}}_{m}\end{cases}$.
10      for $t=T_{m-1}+1$ to $T_{m}$ do
11             Choose arm $I_{t}\sim p_{m}$ where $p_{m}(k)=\widetilde{n}_{k}^{m}/N_{m}$.
@@ 4       Set $\zeta_{m}\leftarrow(m+4)2^{2(m+4)}\ln(K)$, and $\delta_{m}\leftarrow 1/(K\zeta_{m})$
12            Observe the corrupted reward $\widetilde{r}_{t,I_{t}}$ and update the total reward $S_{I_{t}}^{m}=S_{I_{t}}^{m}+\widetilde{r}_{t,I_{t}}$.
13       end for
14
15      Set $r_{k}^{m}\leftarrow\min\{S_{k}^{m}/\widetilde{n}_{k}^{m},1\}$.
16      Set $r_{*}^{m}\leftarrow\top(d)_{k\in[K]}\left\{r_{k}^{m}-\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}\right\}$, where $\top(d)$ represents that the $d$ largest number.
17      Set $\Delta_{k}^{m}\leftarrow\max\{2^{-m},r_{*}^{m}-r_{k}^{m}\}$.
18 end for
Algorithm 4 DS-BARBAT: d-Sets-BARBAT
In this paper, we focus on the $d$-sets setting, which is a special case of semi-bandits where the agent must select $d$ arms in each round. Specifically, let $d\in\{1,2,\ldots,K-1\}$ be a fixed parameter, and define the $d$-sets as
$\mathcal{X}=\left\{\mathbf{x}\in\{0,1\}^{K}\mid\sum_{k=1}^{K}x_{k}=d\right\},$
where $x_{k}=1$ indicates that arm $k$ is selected, and $x_{k}=0$ indicates that arm $k$ is not selected. Following (Zimmert et al., 2019; Dann et al., 2023), let $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{K}$ and $\Delta_{k}=\mu_{k}-\mu_{d}$ for all arms $k>d$.
We extend BARBAT to DS-BARBAT, presented in Algorithm 4. In this setting, the agent observes the rewards of $d$ arms per pull. Similar to MA-BARBAT, we set $N_{m}\leftarrow\lceil K\lambda_{m}2^{2(m-1)}/d\rceil$. However, unlike the traditional MAB where the optimal action is a single arm, here the optimal action is a subset of $d$ arms. Thus, we choose ${\mathcal{K}}_{m}$ to be the $d$ arms with the highest empirical rewards in the previous epoch. When estimating the suboptimality gap, we focus on the $d$ largest empirical values, reflecting the fact that all $d$ selected arms are considered optimal.
The regret bound for DS-BARBAT is as follows, with the proof provided in Appendix C:
Theorem 3
The expected regret of DS-BARBAT satisfies
${\mathbb{E}}\left[R(T)\right]=O\left(dC+\sum_{k=d+1}^{K}\frac{\log(T)\log(KT)}{\Delta_{k}}+\frac{dK\log\left(1/\Delta\right)\log\left(K/\Delta\right)}{\Delta}\right).$
@@
Notice that our DS-BARBAT algorithm can efficient compute the sampling probability $p_{m}$, while FTRL-based methods (Wei and Luo, 2018; Zimmert et al., 2019; Ito, 2021a; Tsuchiya et al., 2023) need to solve a complicated convex optimization problem in each round, which is rather expensive.
4.3 Batched Bandits
1
2Initial round $T_{0}=0$ and $\Delta_{k}^{0}=1$ for all $k\in[K]$.
3for epochs $m=1,2,\cdots$ do
4       Set $a\leftarrow\max\{T^{\frac{1}{2({\mathcal{M}}+1)}},2\}$.
5      Set $\zeta_{m}\leftarrow(m+4)a^{2(m+4)}\ln(aK)$ and $\delta_{m}\leftarrow 1/(K\zeta_{m})$.
6      Set $\lambda_{m}\leftarrow a^{8}\ln{\left(4K/\delta_{m}\right)}$ and $\beta_{m}\leftarrow\delta_{m}/K$.
7      Set $n_{k}^{m}=\lambda_{m}(\Delta_{k}^{m-1})^{-2}$ for all arms $k\in[K]$.
8      Set $N_{m}\leftarrow\lceil K\lambda_{m}a^{2(m-1)}\rceil$ and $T_{m}\leftarrow T_{m-1}+N_{m}$.
9      Select the arm $k_{m}=\mathop{\arg\max}_{k\in[K]}r_{k}^{m-1}$.
10      Set $\widetilde{n}_{k}^{m}=\begin{cases}n_{k}^{m}&k\neq k_{m}\\ N_{m}-\sum_{k\neq k_{m}}n_{k}^{m}&k=k_{m}\end{cases}$.
11      for $t=T_{m-1}+1$ to $T_{m}$ do
12             Choose arm $I_{t}\sim p_{m}$ where $p_{m}(k)=\widetilde{n}_{k}^{m}/N_{m}$.
13            Observe the corrupted reward $\widetilde{r}_{I_{t}}$ and update the total reward $S_{I_{t}}^{m}=S_{I_{t}}^{m}+\widetilde{r}_{I_{t}}$.
14       end for
15
16      Set $r_{k}^{m}\leftarrow\min\{S_{k}^{m}/\widetilde{n}_{k}^{m},1\}$.
17      Set $r_{*}^{m}\leftarrow\max_{k\in[K]}\left\{r_{k}^{m}-\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}\right\}$.
18      Set $\Delta_{k}^{m}\leftarrow\max\{a^{-m},r_{*}^{m}-r_{k}^{m}\}$.
19 end for
Algorithm 5 BB-BARBAT: Batched Bandits-BARBAT
In this setting, the agent can only use data from a batch after it has concluded. The time horizon $T$ is divided into ${\mathcal{M}}$ batches, represented by a grid ${\mathcal{T}}=\{t_{1},t_{2},\dots,t_{{\mathcal{M}}}\}$, where $1\leq t_{1}<t_{2}<\cdots<t_{{\mathcal{M}}}=T$. This grid can be one of two types: 1) Static Grid: The grid is fixed and predetermined before any arm sampling occurs; 2) Adaptive Grid: The value of $t_{j}$ is determined dynamically after observing the rewards up to time $t_{j-1}$, which may incorporate external randomness. Previous works (Perchet et al., 2015; Gao et al., 2019; Zhang et al., 2020; Esfandiari et al., 2021; Jin et al., 2021) have focused on stochastic batched bandits. To the best of our knowledge, no existing works study stochastic batched bandit problem under adversarial corruptions.
We extend BARBAT to the batched bandit setting by introducing BB-BARBAT, presented in Algorithm 5. Building upon BARBAT’s epoch-based structure, we increase the epoch length $N_{m}$ as we progress through the batches. We introduce a parameter $a\leftarrow\max\{T^{\frac{1}{2({\mathcal{M}}+1)}},2\}$, ensuring $\sum_{m=1}^{M-1}N_{m}<T$ and $\sum_{m=1}^{M-1}N_{m}\geq T$, where $M$ is the number of epochs. This structure ensures that if ${\mathcal{M}}\leq\log(T)$, BB-BARBAT runs for exactly ${\mathcal{M}}$ epochs. If ${\mathcal{M}}>\log(T)$, then we have $a=2$, leading to $\log(T)$ epochs in BB-BARBAT to maintain robustness. In this case, BB-BARBAT is exactly the same as BARBAT. We show that BB-BARBAT achieves the following regret bound, with the proof provided in Appendix D:
Theorem 4
For an ${\mathcal{M}}$-batched, $K$-armed stochastic bandit problem with time horizon $T$, where the number of epochs satisfies $M=\min\{{\mathcal{M}},\log(T)\}$, the expected regret of BB-BARBAT is
${\mathbb{E}}\left[R(T)\right]=O\left(CT^{\frac{1}{M+3}}+T^{\frac{4}{M+3}}\left(\sum_{\Delta_{k}>0}\frac{M\log(KT)}{\Delta_{k}}+\frac{K\log(T)\log(1/\Delta)\log(K/\Delta)}{M\Delta}\right)\right).$
We provide the lower bound of corrupted batched bandits in the following theorem.
Theorem 5
For an ${\mathcal{M}}$-batched, $K$-armed stochastic bandit problem with time horizon $T$, corruption level $C$ and any static grids, the regrets can be lower bounded as
$R(T)\geq\Omega\left(T^{\frac{1}{{\mathcal{M}}}}\left(K+C^{1-\frac{1}{{\mathcal{M}}}}\right)\right).$
Remark 3
Our upper bound nearly matches the proposed lower bound when ${\mathcal{M}}\geq\log(T)$. However, there is still a gap between our upper bound and the lower bound. We think that our lower bound is nearly tight, and how to achieve optimal regret upper bounds for corrupted batched bandits is an interesting open question.
4.4 Multi-Agent Multi-Armed Bandits
In CMA2B, each agent can broadcast messages in one round to all $V$ agents. For simplicity, we define the total communication cost as:
[…]
and we aim to keep communication costs sub-linear in $T$.
We extend BARBAT to MA-BARBAT (Algorithm 6) with only minor adjustments. In MA-BARBAT, we split the epoch into $V$ parts, allowing each agent to pull arms with identical probabilities and then broadcast the total rewards at the end. This ensures the combined reward is equivalent to that in a (unsplit) epoch by a single agent. Each agent $v$ broadcasts $\{S_{v,k}^{m}\}_{k=1}^{K}$ at the end of an epoch, enabling all agents to compute $\Delta_{k}^{m-1}$ from others’ observations.
1
2
3Initialization: Set the initial round $T_{0}=0$, $\Delta_{k}^{0}=1$, and $r_{k}^{0}=0$ for all $k\in[K]$.
4for epoch $m=1,2,\dots$ do
5       Each agent sets $\zeta_{m}\leftarrow(m+4)2^{2(m+4)}\ln(VK)$ and $\delta_{m}\leftarrow 1/VK\zeta_{m}$.
6      Each agent sets $\lambda_{m}\leftarrow 2^{8}\ln\left(4K/\delta_{m}\right)/V$, $\beta_{m}\leftarrow\delta_{m}/VK$.
7      Each agent sets $n_{k}^{m}\leftarrow\lambda_{m}(\Delta_{k}^{m-1})^{-2}$ for all $k\in[K]$.
8      Each agent sets $N_{m}\leftarrow\lceil K\lambda_{m}2^{2(m-1)}\rceil$ and $T_{m}\leftarrow T_{m-1}+N_{m}$.
9      Each agent selects the arm $k_{m}=\mathop{\arg\max}_{k\in[K]}r_{k}^{m-1}$. In case of ties (multiple arms with the same $r_{k}^{m-1}$), select the arm with the smallest index as $k_{m}$.
10      Each agent sets:
$\widetilde{n}_{k}^{m}=\begin{cases}n_{k}^{m}&\text{if }k\neq k_{m}\\ N_{m}-\sum_{k\neq k_{m}}n_{k}^{m}&\text{if }k=k_{m}\end{cases}$
11      for $t=T_{m-1}+1$ to $T_{m}$ do
12             Each agent chooses arm $I_{t}\sim p_{m}$, where $p_{m}(k)=\widetilde{n}_{k}^{m}/N_{m}$.
13            Each agent observes the corrupted reward $\widetilde{r}_{t,I_{t}}$ and updates the total reward: $S_{I_{t}}^{m}=S_{I_{t}}^{m}+\widetilde{r}_{t,I_{t}}.$
14       end for
15
16      Each agent $v$ broadcasts the messages $\{S_{v,k}^{m}\}_{k\in[K]}$ to other agents.
17      Each agent sets $r_{k}^{m}\leftarrow\min\{\sum_{v\in[V]}S_{v,k}^{m}/(V\widetilde{n}_{k}^{m}),1\}$.
18      Each agent sets $r_{*}^{m}\leftarrow\max_{k\in[K]}\left\{r_{k}^{m}-\sqrt{\frac{4\ln(4/\beta_{m})}{V\widetilde{n}_{k}^{m}}}\right\}$.
19      Each agent set $\Delta_{k}^{m}\leftarrow\max\{2^{-m},r_{*}^{m}-r_{k}^{m}\}$.
20 end for
Algorithm 6 MA-BARBAT: Multi-Agent-BARBAT
We show that our MA-BARBAT algorithm have the following regret bound. The proof is given in Appendix E.
Theorem 6
With only $V\log(VT)$ communication cost, we can guarantee the expected individual regret of each agent $i$ as follows:
${\mathbb{E}}\left[R(T)\right]=O\left(\frac{C}{V}+\sum_{\Delta_{k}>0}\frac{\log(VT)\log(KVT)}{V\Delta_{k}}+\frac{K\log\left(1/\Delta\right)\log\left(KV/\Delta\right)}{V\Delta}\right).$
Compared with previous works (Liu et al., 2021) which achieves $O(VC+\frac{K\log^{2}(T)}{\Delta})$ regret and the concurrent work (Ghaffari et al., 2024) which achieve $O(C/V+\frac{K\log^{2}(T)}{V\Delta})$, our regret bound is better.
Remark 4
Our regret bound shows that collaboration can reduce the individual regret of each agent by a factor of $V$. Notice that some previous works Wang et al. (2022) for CMA2B without corruptions only require $o(\log(T))$ communication cost, but these methods significantly depends on the assumption of unique best arm. For the case with multiple best arms, it is still unknown whether $o(\log(T))$ communication cost is enough.
@@ 5 Conclusion
Building on the work of Gupta et al. (2019), we propose the BARBAT framework by elaborately choosing epoch lengths $N_{m}$ and the parameters $\delta_{m}$. we also extend BARBAT to various corrupted bandit settings including strongly observable graph bandits, $d$-set semi-bandits, multi-agent bandits and batched bandits. The analysis shows that our algorithms can achieve near-optimal regret bounds for the adversarial corruption environments. Compared with FTRL-based methods, BARBAT and its extensions are more efficient and do not require the assumption of unique optimal action. Also, FTRL-based methods are intractable to parallelize so that it is not applicable to the multi-agent bandits and batched bandits. These advantages make the BARBAT framework an appealing alternative to FTRL algorithms for many corrupted bandits settings.
One future work is to improve the $\log^{2}(T)$ dependence in our regret bound to $\log(T)$. Investigating extensions in other bandits settings such as linear bandits is also a interesting future direction.
Appendix A Proof of Theorem 1
A.1 Notations

### v2

@@ 4.4 $d$-set Semi-bandits
Algorithm 5 DS-BARBAT: d-Set-BARBAT
1:Initialization: Set the initial round $T_{0}\leftarrow 0$, $\Delta_{k}^{0}\leftarrow 1$, and $r_{k}^{0}\leftarrow 0$ for all $k\in[K]$.
2:for epochs $m=1,2,\cdots$ do
3:  Set $\zeta_{m}\leftarrow(m+4)2^{2(m+4)}\ln(K)$, and $\delta_{m}\leftarrow 1/(K\zeta_{m})$
4:  Set $\lambda_{m}\leftarrow 2^{8}\ln{\left(4K/\delta_{m}\right)}$ and $\beta_{m}\leftarrow\delta_{m}/K$.
5:  Set $n_{k}^{m}\leftarrow\lambda_{m}(\Delta_{k}^{m-1})^{-2}$ for all arms $k\in[K]$.
6:  Set $N_{m}\leftarrow\lceil K\lambda_{m}2^{2(m-1)}/d\rceil$ and $T_{m}\leftarrow T_{m-1}+N_{m}$.
7:  Set the arm subsets ${\mathcal{K}}_{m}\leftarrow\operatorname*{arg\,max}_{k\in[K]}r_{k}^{m-1}$ with $|{\mathcal{K}}_{m}|=d$.
8:  Set $\widetilde{n}_{k}^{m}=\begin{cases}n_{k}^{m}&k\not\in{\mathcal{K}}_{m}\\ N_{m}-\sum_{k\not\in{\mathcal{K}}_{m}}n_{k}^{m}/d&k\in{\mathcal{K}}_{m}\end{cases}$.
9:  for $t=T_{m-1}+1$ to $T_{m}$ do
10:   Choose action $\mathbf{x}_{t}\sim p_{m}$ where $p_{m}(k)=\widetilde{n}_{k}^{m}/N_{m}$.
@@
11:   For each arm $k$ with $\mathbf{x}_{t}(k)=1$, observe $\widetilde{r}_{t,k}$ and update the total reward $S_{k}^{m}\leftarrow S_{k}^{m}+\widetilde{r}_{t,k}$.
12:  end for
13:  Set $r_{k}^{m}\leftarrow\min\{S_{k}^{m}/\widetilde{n}_{k}^{m},1\}$.
14:  Set $r_{*}^{m}\leftarrow\top_{d}\Big(\big\{r_{k}^{m}-\sqrt{\tfrac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}\big\}_{k\in[K]}\Big)$, where $\top_{d}({\mathcal{A}})$ returns the $d$-th largest value in $\mathcal{A}$}.
15:  Set $\Delta_{k}^{m}\leftarrow\max\{2^{-m},r_{*}^{m}-r_{k}^{m}\}$.
16:end for
We extend BARBAT to the $d$-set semi-bandit setting and propose the DS-BARBAT (d-Set BARBAT) algorithm in Algorithm 5. In this setting, the optimal action is a subset of $d$ distinct arms. Thus, DS-BARBAT estimates the best action ${\mathcal{K}}_{m}$ as set of $d$ distinct arms with the highest empirical rewards in the previous epoch. When estimating the suboptimality gap, DS-BARBAT considers the arm with the $d$-th largest empirical reward, since the set of the first $d$ arms is the optimal action. The regret bound for DS-BARBAT is presented as follows, with the proof provided in Appendix B.5:
Theorem 6.
The regret of DS-BARBAT satisfies
$R(T)=O\left(dC+\sum_{k=d+1}^{K}\frac{\log(T)\log(KT)}{\Delta_{k}}+\frac{dK\log\left(1/\Delta\right)\log\left(K/\Delta\right)}{\Delta}\right).$
Remark 9.
@@
Notice that our DS-BARBAT algorithm can efficiently compute the sampling probability $p_{m}$, while FTRL-based methods (wei2018more; zimmert2019beating; ito2021hybrid; tsuchiya2023further) need to solve a complicated convex optimization problem in each round, which is rather expensive. The time comparison in Table 3 in Appendix C.3 demonstrates that our DS-BARBAT algorithm is significantly faster than these FTRL-based methods.
5 Experiments
Refer to caption Refer to caption Refer to caption Refer to caption
(a) K = 12, C = 2000 (b) K = 12, C = 5000 (c) K = 16, C = 2000 (d) K = 16, C = 5000
Figure 1: Comparison between MA-BARBAT, DRAA, IND-BARBAR and IND-FTRL in cooperative multi-agent multi-armed bandits.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) K = 12, C = 2000 (b) K = 12, C = 5000 (c) K = 16, C = 2000 (d) K = 16, C = 5000
Figure 2: Comparison between SOG-BARBAT, Shannon-FTRL, and Tsallis-FTRL in strongly observable graph bandits.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) K = 12, C = 2000 (b) K = 12, C = 5000 (c) K = 16, C = 2000 (d) K = 16, C = 5000
Figure 3: Comparison between DS-BARBAT, HYBRID, LBINF, LBINF_LS and LBINF_GD in $d$-set semi-bandits.
In this section, we present numerical results to demonstrate the robustness and effectiveness of our framework. All experiments are implemented in Python 3.11 and conducted on a Windows laptop equipped with 16 GB of memory and an Intel i7-13700H processor. For all experiments, we run $50$ independent trials, report the average results and use shaded regions to represent standard deviations. The detailed experimental settings and time comparisons are deferred to Appendix C.
5.1 Cooperative Multi-agent Multi-Armed Bandits
We compare our framework with DRAA ghaffari2025multi, IND-BARBAR gupta2019better, and IND-FTRL zimmert2021tsallis in a multi-agent setting with $V=10$. Here, IND-BARBAR and IND-FTRL serve as non-cooperative baselines, where each agent runs the algorithm locally without any communication. We do not compare with liu2012learning since the baseline DRAA is an improved version of their method. We present the experimental results in Figure 1, which shows that our BARBAT framework outperforms all baseline methods, reflecting the advantage of collaboration and verifying our theoretical analysis. In addition, when $C$ is increased, the increase in individual regret of our framework is the smallest, which shows that the robustness of our framework is stronger than that of all baseline methods.
5.2 Strongly Observable Graph Bandits
We compare our SOG-BARBAT algorithm with Shannon-FTRL ito2022nearly and Tsallis-FTRL dann2023blackbox using two directed graphs as illustrated in Figure 4 in Appendix C.2. We do not include the Elise lu2021stochastic algorithm in the experiment since it only works for undirected graphs. The experimental results are presented in Figure 2, showing that our SOG-BARBAT algorithm significantly outperforms all baseline methods and verifying the robustness of our method.
5.3 $d$-Set Semi-bandits
In $d$-set semi-bandits, we set $d=3$ for $K=12$, and $d=4$ for $K=16$. We compare our framework with HYBRID zimmert2019beating, LBINF ito2021hybrid, LBINF_LS tsuchiya2023further and LBINF_GD tsuchiya2023further. As shown in Figure 3, our DS-BARBAT algorithm significantly outperforms all baseline methods.
@@ 6 Conclusion
Building on the BARBAR algorithm gupta2019better, we propose the BARBAT framework by elaborately designing the epoch lengths $N_{m}$ and the parameters $\delta_{m}$. We further extend BARBAT to various corrupted bandit settings, including multi-agent bandits, batched bandits, strongly observable graph bandits and $d$-set semi-bandits. Our theoretical analysis shows that these algorithms achieve near-optimal regret bounds for adversarially corrupted environments. Compared with FTRL-based approaches, BARBAT and its extensions are more computationally efficient and do not rely on the assumption of a unique optimal action. Moreover, since FTRL-based methods are difficult to parallelize, they are unsuitable for multi-agent and batched settings. These advantages make the BARBAT framework an attractive alternative to FTRL-based algorithms across a wide range of corrupted bandit problems. One future direction is to improve the $\log^{2}(T)$ dependence in our regret bound to $\log(T)$. Investigating extensions in other bandit settings such as linear bandits is also an interesting future direction.
Acknowledgement
This work was supported by Natural Science Foundation of China (No. 62306116), and in part by the Key Program of National Natural Science Foundation of China (No. 62432007).
Appendix A OODS algorithm for strongly observable graph bandits
In this section, we present the OODS algorithm for strongly observable graph bandits. For any graph $G=(V,E)$, the time complexity of the algorithm is $O(|V|^{2}|E|)$.
Algorithm 6 OODS: Obtain an out-domination Set
1:Input: a graph $G$.
2:Output: an out-domination set $D$.
3:while the graph $G$ is not empty do
4:  if the graph $G$ is an acyclic graph then
5:   Find all no-root vertices and add them to the set $D$.
6:   Remove all no-root vertices and their out-degree neighbors from the graph $G$.
7:  else
8:   Select the vertex with the largest out-degree in the graph $G$ and add it to the set $D$.
9:   Remove the vertex and its out-degree neighbors from the graph $G$.
10:  end if
11:end while
A “no-root vertex” is defined as a vertex with no incoming edges other than its self-loop. To ensure that the rewards for no-root vertices are observed, OODS pulls these vertices directly. However, in some graphs, there may be no such vertices. In such cases, OODS applies a greedy strategy: it selects the vertex with the highest out-degree and adds it to $D$. The procedure proceeds as follows: if a cycle is detected in $G$, the greedy strategy is employed; otherwise, all no-root vertices are added to $D$. The vertices in $D$ and their out-degree neighbors are then removed from the graph $G$, and the process repeats until $G$ is empty, ensuring that $D$ forms an out-domination set covering $G$.
At a high level, the OODS algorithm finds an out-domination set $D^{m}$ such that pulling the arms $k\in D^{m}$ allows us to observe the rewards of all arms. The loop iteratively removes arms once they meet the observation requirement. After each removal, the independence number of the remaining graph may decrease, so the required dominating set becomes smaller—thereby reducing the number of pulls allocated to suboptimal arms.
Appendix B Proof Details
B.1 Proof of Theorem 1
B.1.1 Notations